OpenStudy (anonymous):

Can someone check my work? I did an assignment for algebra 2 & I want to make sure that it's right.

OpenStudy (anonymous):

1. Write a word problem describing the time it takes to complete an activity individually and with a friend. For example, if John takes 2 hours to mow his lawn and it takes his sister Maria 4 hours to mow the same lawn, how long would it take John and Maria to mow the lawn together? 2. Write a rational equation based upon the word problem you created. The rational equation based upon the scenario above would be 1 over 2 + 1 over 4 = 1 over x. Each fraction represents the amount of the lawn mowed in one hour. The fraction of 1 over 2 is John’s portion. Maria’s share is represented by 1 over 4. The time it would take for both of them to mow the lawn is represented by 1 over x. 3. Solve the rational equation. Show your work. Here's the work for those questions.. 1. It takes Joe 7 hours to paint a house & Bill 5 hours. How long would it take them to paint a house together? 2. 1/7 + 1/5 = 1/x. Each fraction represents how much of the house they painted in an hour. Joe is the 1/7 and bill is the 1/5 and the 1/x is what they painted together. 3. 1/5 + 1/7 = 1/x 1/5 + 1/7 = 5 + 7/35 = 12/35 12/35 = 1/x 12/35x = 1 12x=35 x = 2.91 or 2.916 2.916 rounds to 2.92 but there are 60 minutes in an hour so u gotta take out the 60 which makes another hour making it 3 and leaving 32 minutes still. So all in all it takes them 3 hours & 32 minutes to do it together.